OpenStudy (anonymous):

h-(-15)=1.5 (with steps) h-(-15)=1.5 (with steps) @Mathematics

OpenStudy (anonymous):

h-(-15) =1.5 h + 15 = 1.5 h = 1.5 - 15 h = -13.5